Unit 1 Review:

5) What is domain of

f(x) = { 3x + 2, -3 < x < 1

-2x + 5, 1 < x < 4


6) What is range of

f(x) = { 3x+2, -3 < x < 1

-2x + 5, 1 < x < 4


1
Expert's answer
2021-09-23T14:17:50-0400

Given that: "f(x) = \begin{cases}
 3x+2 &\text{if } -3<x<1 \\
 -2x+5 &\text{if } 1<x<4
\end{cases}"

(1) From the given data, we can say directly that the domain of the function is: "x\in(-3,1)\cup(1,4)"

(2) Range:

At "x=-3,f(-3)=-9+2=-7\\"

At "x=1, f(1)=3+2=5"

At "x=4, f(4)=-8+5=-3"

So, range of "f(x)" is: "(-7,5)"